what are the prime factors of 80,60,40?
Answer by radh(108)

Algebra.com has solvers to find the factors of any number.

Solved by pluggable solver: Find factors of any number
80 is NOT a prime number: 80 = 2 * 2 * 2 * 2 * 5

Work Shown

80 is divisible by 2: 80 = 40 * 2.
40 is divisible by 2: 40 = 20 * 2.
20 is divisible by 2: 20 = 10 * 2.
10 is divisible by 2: 10 = 5 * 2.
5 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.

You can see from the work shown, that obviously 5 is a prime factor, and so is 2 because 2 is not divisible by anything but itself and 1.

So, to find factors, you can use the solver (or factor it using logic), and look for the numbers that are not divisible by anything other than itself and 1. Try it for the other numbers.